# GL1-02-06-24-2-supplement

### v2

@@ 5 Underreported Outcomes
It should be emphasized that the distributional parallel trends solely set constraints on the change of the marginal distributions of counterfactual outcomes over time. However, these assumptions do not place any constraints on the evolution of the dependence structure between the outcome and reporting decisions. Additionally, it is noteworthy that the identification result presented in Proposition 4 requires knowledge of neither the marginal distributions nor the Copula function, thereby classifying it as semiparametric. Nevertheless, when transitioning to estimation, one has to select specific parametric forms for the Copula function $\overline{C}_{gt}(\cdot,\cdot)$ as well as the distributions $F_{-\varepsilon_{gt}}$ and $F_{-\nu_{gt}}$. This is because the identification based on the real analytic assumption is not “constructive” and does not provide a direct pathway to construct estimators. As discussed earlier, however, due to the partially observable nature of the consumption and reporting decisions, it appears difficult to bypass this real analytic assumption without relying on excluded variables with large support.
Remark 1: Note that, under the assumption of index sufficiency and instrument validity, the above assumption setup not only allows for specification tests of the functional form, but also for the presence of misreporting. Specifically, starting with the test of correct functional form and focusing on a specific choice of $(g,t)\neq(1,1)$, let $M\left(j;\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{Z}_{gt}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right):=\Pr\left(C_{gt}\geq j\mid\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{Z}_{gt}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)$ denote the nonparametric tail probability of observed consumption at level $j\in\{0,1,2\}$, where $\widetilde{\boldsymbol{\eta}}_{gt}$ and $\widetilde{\boldsymbol{\pi}}_{gt}$ are the slope coefficient vectors normalized by $\|\boldsymbol{\eta}_{gt}\|$ and $\|\boldsymbol{\pi}_{gt}\|$, respectively. Note that this tail probability is nonparametrically identified under Assumption 7 and index sufficiency, see for instance Donkers and Schafgans (2008) or Ahn et al. (2018). Then, under Assumptions 6 and 7, we may construct a test for the null hypothesis:
$H_{0}^{C}:\overline{C}_{gt}\left(F_{-{\varepsilon}_{gt}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{gt}-\kappa_{j}}{\lambda_{gt}}\right),F_{-{\nu}_{gt}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}_{gt}-\iota_{j}}{\zeta_{gt}}\right)\right)=M\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{z}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)$
for all $j\in\{0,1,2\}$ and $(\mathbf{x},\mathbf{z})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}$, where $\mathcal{X}_{gt}$ and $\mathcal{Z}_{gt}$ denote compact subsets of the support of $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$, respectively, vs.
$H_{A}^{C}:\overline{C}_{gt}\left(F_{-{\varepsilon}_{gt}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{gt}-\kappa_{j}}{\lambda_{gt}}\right),F_{-{\nu}_{gt}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}_{gt}-\iota_{j}}{\zeta_{gt}}\right)\right)\neq M\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{z}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)$
for at least some $j\in\{0,1,2\}$ and $(\mathbf{x},\mathbf{z})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}$. Similarly, we may also use the fact that $M\left(j;\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{Z}_{gt}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)$ is nonparametrically identified under our assumptions to test for the presence of misreporting in our semiparametric model setup. That is, given index sufficiency and Assumption 7, we may test:
$H_{0}^{R}:M\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{z}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)=\overline{M}\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt}\right)$
for all $j\in\{0,1,2\}$ and all $(\mathbf{x},\mathbf{z})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}$ vs.
$H_{0}^{R}:M\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{z}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)\neq\overline{M}\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt}\right)$
for at least some $j\in\{0,1,2\}$ and $(\mathbf{x},\mathbf{z})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}$, where $\overline{M}\left(j;\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt}\right):=\Pr\left(C_{gt}\geq j\mid\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt}\right)$. We comment on suitable test statistics for the above tests in Remark 3 in Section 7.2.
6 Distributional Treatment Effects
@@
To identify the DTTs, $\tau_{11}^{c}(j)$, $j\in\{0,1,2\}$, as outlined in (2), we start by considering the scenario where $Y_{gt}(d)=C_{gt}(d)$ almost surely, for $d\in\{0,1\}$. Although modeling $Y_{11}(1)$ is per se not essential in DiD settings, it becomes indispensable when addressing scenarios involving misreporting. Therefore, to simplify the notation and facilitate the transition to the model with misreporting, we also postulate an analogous model structure for $Y_{11}(1)$ as for $Y_{11}(0)$ without misreporting. Specifically, drawing on the model for $Y_{11}^{\ast}(0)$ in (13), we define the latent consumption in the presence of treatment as:
@@
where we introduced the superscript (1) to distinguish the parameters of the potential outcome under treatment from those under control. From this point forward, we will accordingly add the superscript $(0)$ to parameters in the absence of treatment. Given the similarity in the structure of $Y_{11}^{\ast}(1)$, we adopt assumptions similar to those for $Y_{11}^{\ast}(0)$ in Section 4.1 above:
@@
$\varepsilon_{11}^{(1)}$ is strictly stationary within each group with zero median and unit variance, independent of $\mathbf{X}_{11}$, and has an unknown, continuous, and strictly monotonic CDF $F_{\varepsilon_{11}^{(1)}}$. Besides, $\eta_{1,11}^{(1)}\neq 0$.
We readily obtain the following proposition for identification of the conditional DTTs for consumption, in the absence of misreporting:
Proposition 5.
Suppose Assumptions 1-5 and 8 hold, and consider the model in (13) and (18). Also, suppose distributional parallel trends hold as in Definition 1 for $Y_{gt}^{*}(0)$. Then, $\tau_{11}^{c}(j\mid{\mathbf{x}})$ is identified for all $j\in\{0,1,2\}$ and $\mathbf{x}\in\mathcal{X}_{11}$ as:
$\tau_{11}^{c}(j\mid{\mathbf{x}})=$ $\left(F_{\varepsilon_{11}^{(1)}}\left(\frac{\kappa_{j+1}-\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}}{\lambda^{(1)}_{11}}\right)-F_{\varepsilon_{11}^{(1)}}\left(\frac{\kappa_{j}-\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}}{\lambda^{(1)}_{11}}\right)\right)$
$-\left(F_{\varepsilon_{11}^{(0)}}\left(\frac{\kappa_{j+1}-\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(0)}_{11}}{\lambda_{11}^{(0)}}\right)-F_{\varepsilon_{11}^{(0)}}\left(\frac{\kappa_{j}-\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(0)}}{\lambda_{11}^{(0)}}\right)\right).$
Next, we turn to the case where the outcome is subject to underreporting. Drawing on the model for $R_{11}^{\ast}(0)$ in (14), We define:
$R_{11}^{\ast}(1)=\overline{\mathbf{Z}}_{11}^{\prime}\overline{\boldsymbol{\pi}}_{11}^{(1)}+\zeta_{11}^{(1)}\nu_{11}^{(1)}.$ (19)
We adopt assumptions similar to those for $Y_{11}^{\ast}(0)$ in Section 4.2 above:
Assumption 9.
$-\varepsilon_{11}^{(1)}$ and $-\nu_{11}^{(1)}$ are strictly stationary within each group with zero median and unit variance, jointly independent of $(\mathbf{X}_{11}^{\prime},\mathbf{Z}_{11}^{\prime})^{\prime}$, and have unknown, real analytic and strictly monotonic CDFs $F_{-\varepsilon_{11}^{(1)}}$ and $F_{-\nu_{11}^{(1)}}$, respectively. Besides, $\eta_{1,11}^{(1)}\neq 0$ and $\pi_{1,11}^{(1)}\neq 0$ and there exists some Cartesian product of two open intervals $I_{1}\times I_{2}\subseteq\mathbb{R}\times\mathbb{R}$ such that
$\operatorname{Pr}\left(\left(X_{1,11},Z_{1,11}\right)\in I_{1}\times I_{2}\mid\overline{\mathbf{W}}_{11}\right)>0$
with probability one, where $\overline{\mathbf{W}}_{11}$ is defined as in Assumption 7. Also, the copula function $\overline{C}_{11}^{(1)}(\cdot,\cdot)$ defined below is strictly monotonic and real analytic in each of its coordinates:
@@
Suppose Assumptions 6-7 and 9 hold, and consider the model in (13)-(14) and (18)-(19). Also, suppose distributional parallel trends hold as in Definition 1 for $Y_{gt}^{*}(0)$ as well as $R_{gt}^{*}(0)$. Then, $\tau_{11}^{c}(j)$ and $\tau_{11}^{r}(j)$ are identified for all $j\in\{0,1,2\}$, $\mathbf{x}\in\mathcal{X}_{11}$ and $\mathbf{z}\in\mathcal{Z}_{11}$ as:
$\tau_{11}^{c}(j\mid{\mathbf{x}})=$ $\left(F_{-\varepsilon_{11}^{(1)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}-\kappa_{j+1}}{\lambda^{(1)}_{11}}\right)-F_{-\varepsilon_{11}^{(1)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}-\kappa_{j}}{\lambda^{(1)}_{11}}\right)\right)$
@@
$\tau_{11}^{r}(j\mid{\mathbf{z}})=$ $\left(F_{-\nu_{11}^{(1)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(1)}_{11}-\iota_{j+1}}{\zeta^{(1)}_{11}}\right)-F_{-\nu_{11}^{(1)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(1)}_{11}-\iota_{j}}{\zeta^{(1)}_{11}}\right)\right)$
@@
7 Estimation
In this section, we discuss the estimation of the DTTs $\tau_{11}^{c}(j)$. For each observational unit $i$, let $G_{i}$ and $T_{i}$ denote binary indicators of group membership (with $G_{i}=1$ indicating being in the treated group) and time period (with $T_{i}=1$ indicating being after the treatment), respectively. Following Sant’Anna and Zhao (2020), we write the observed outcomes and the covariates as $(C_{it},\mathbf{X}_{it}^{\prime},\mathbf{Z}_{it}^{\prime})^{\prime}$, where the subscript $t$ denotes the time period. This allows to accommodate various sampling frameworks, including repeated cross-section sampling, where observations are drawn conditional on a specific group $G_{i}=g$ and time period $T_{i}=t$, and panel sampling that draws observation pairs conditional on $G_{i}=g$, see Assumption 10 below. Consequently, we define $C_{i}=C_{i1}T_{i}+C_{i0}(1-T_{i})$ and $Y_{i}=Y_{i1}T_{i}+Y_{i0}(1-T_{i})$, with regressors defined in a similar manner. It is important to recall that in the absence of “false zeros”, we observe $Y_{it}=C_{it}$ almost surely.
7.1 Estimation Without Misreporting
We start with the estimation of the counterfactual distribution in the absence of misreporting, as described in Section 4, focusing initially on a parametric model without covariates detailed in Section 4.1. Recalling that $\kappa_{0}=-\infty$, $\kappa_{1}=0$, $\kappa_{2}=1$, $\kappa_{3}=+\infty$, we consider a sample $\{C_{i},G_{i},T_{i}\}_{i=1}^{n}$ with the properties specified in Assumption 10 below. Also, suppose Assumptions 1 and 8 hold and consider the models in (12) and (18) (no covariates for the latter), where the distributions of $\varepsilon_{gt}^{(d)}$ are known for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\}$. To avoid notational complexity, we assume that the error term distributions are invariant across groups, time and treatment statuses, i.e., $F_{\varepsilon_{gt}^{(d)}}=F_{\varepsilon}$. We define $\boldsymbol{\theta}^{(d)}_{gt}:=\left(\eta^{(d)}_{gt},\lambda^{(d)}_{gt}\right)^{\prime}$ and aggregate these into $\boldsymbol{\theta}:=\left(\boldsymbol{\theta}^{(0)^{\prime}}_{00},\boldsymbol{\theta}^{(0)^{\prime}}_{01},\boldsymbol{\theta}^{(0)^{\prime}}_{10},\boldsymbol{\theta}^{(1)^{\prime}}_{11}\right)^{\prime}$. The ML estimator of $\boldsymbol{\theta}$ is formulated as follows:
$\widehat{\boldsymbol{\theta}}:=$ $\arg\min_{\boldsymbol{\theta}\in\mathcal{B}}\sum_{i=1}^{n}\sum_{j=0}^{2}\mathbb{I}\left[Y_{i}=j\right]\left[\sum_{(g,t)\neq(1,1)}\mathbb{I}\left[G_{i}=g,T_{i}=t\right]\ln\left(F_{\varepsilon}\left(\frac{\kappa_{j+1}-\eta_{gt}^{(0)}}{\lambda_{gt}^{(0)}}\right)-F_{\varepsilon}\left(\frac{\kappa_{j}-\eta_{gt}^{(0)}}{\lambda_{gt}^{(0)}}\right)\right)\right.$
$+\left.\mathbb{I}\left[G_{i}=1,T_{i}=1\right]\ln\left(F_{\varepsilon}\left(\frac{\kappa_{j+1}-\eta^{(1)}_{11}}{\lambda^{(1)}_{11}}\right)-F_{\varepsilon}\left(\frac{\kappa_{j}-\eta^{(1)}_{11}}{\lambda^{(1)}_{11}}\right)\right)\right].$ (20)
where $\mathcal{B}$ denotes the parameter space, $\mathbb{I}\left[A\right]$ is the indicator function that equals 1 when $A$ is true and 0 otherwise.
This step allows for the construction of DTT estimates, $\tau_{11}^{c}(j)$ for $j\in\{0,1,2\}$:
@@ 7.1 Estimation Without Misreporting
$\widehat{\tau}^{c}_{11}(j):=\left[F_{\varepsilon}\left(\frac{\kappa_{j+1}-\widehat{\eta}^{(1)}_{11}}{\widehat{\lambda}^{(1)}_{11}}\right)-F_{\varepsilon}\left(\frac{\kappa_{j}-\widehat{\eta}^{(1)}_{11}}{\widehat{\lambda}^{(1)}_{11}}\right)\right]-\left[F_{\varepsilon}\left(\frac{\kappa_{j+1}-\widehat{\eta}^{(0)}_{11}}{\widehat{\lambda}^{(0)}_{11}}\right)-F_{\varepsilon}\left(\frac{\kappa_{j}-\widehat{\eta}^{(0)}_{11}}{\widehat{\lambda}^{(0)}_{11}}\right)\right]$ (21)
where
$\widehat{\eta}^{(0)}_{11}:=\widehat{\eta}^{(0)}_{01}+\frac{\widehat{\eta}^{(0)}_{10}-\widehat{\eta}^{(0)}_{00}}{\widehat{\lambda}^{(0)}_{00}/\widehat{\lambda}^{(0)}_{01}}\quad\text{and}\quad\widehat{\lambda}^{(0)}_{11}:=\frac{\widehat{\lambda}^{(0)}_{01}\widehat{\lambda}^{(0)}_{10}}{\widehat{\lambda}^{(0)}_{00}}.$
For the inference on the treatment parameters $\tau_{11}^{c}(j)$, it is crucial to derive the asymptotic distribution of the ML estimator. To do so, we impose the following assumptions:
Assumption 10.
It holds that either (i) conditional on $G_{i}=g$ and $T_{i}=t$ with $g,t\in\{0,1\}$, $\{C_{i}\}_{i=1}^{n}$ is an i.i.d. random sample from the corresponding subpopulation or (ii) conditional on $G_{i}=g$ with $g\in\{0,1\}$, $\{C_{i0},C_{i1}\}_{i=1}^{n}$ is an i.i.d. random sample from the corresponding subpopulation, either $g=0$ or $g=1$.
Assumption 11.
The ML estimator of $\boldsymbol{\theta}$ defined in (7.1) satisfies a Central Limit Theorem:
$\sqrt{n}\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(\boldsymbol{0},\boldsymbol{\Sigma}\right),$
where $\boldsymbol{\Sigma}$ is a positive definite matrix.
The first part (i) of Assumption 10 addresses the case where repeated cross-section data is available, while part (ii) covers the case of panel data. While Assumption 10 only accommodates the “conditional” sampling scheme where group membership and time are fixed, the assumption could readily be extended to allow for other schemes where $G_{i}$ and $T_{i}$ are random (see Sant’Anna and Zhao, 2020). Assumption 11 on the other hand is a high-level condition, valid under the stated model assumptions as well as certain primitive conditions on the Hessian and the population shares of each group. In particular, it implies that $n_{gt}/n\rightarrow K$ (with $0<K<\infty$) as $n\rightarrow\infty$ for each $g\in\{0,1\}$, $t\in\{0,1\}$, where $n_{gt}$ denotes the number of observations in the group $g$ and time period $t$.
Proposition 7.
Suppose Assumptions 10-11 hold. Then, for each $j\in\{0,1,2\}$:
$\sqrt{n}\left(\widehat{\tau}^{c}_{11}(j)-\tau_{11}^{c}(j)\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(\boldsymbol{0},\frac{\partial\tau_{11}^{c}(j)}{\partial{{\boldsymbol{\theta}}}}^{\prime}\boldsymbol{\Sigma}\frac{\partial\tau_{11}^{c}(j)}{\partial{{\boldsymbol{\theta}}}}\right),$
where the definition of $\frac{\partial\tau_{11}^{c}(j)}{\partial{{\boldsymbol{\theta}}}}$ is given in Section LABEL:sec:derivatives of the supplementary material.
The components of the asymptotic variance can be readily estimated. Specifically, the asymptotic variance matrix $\boldsymbol{\Sigma}$ can be replaced by a standard heteroskedasticity robust estimator, denoted as $\widehat{\boldsymbol{\Sigma}}_{R}$, while the derivatives $\frac{\partial\tau_{11}^{c}(j)}{\partial{{\boldsymbol{\theta}}}}$ can be substituted with a consistent estimator $\widehat{\partial\tau_{11}^{c}(j)/\partial{{\boldsymbol{\theta}}}}$.
Remark 2: Note that the identification argument for the semiparametric model from Subsection 4.2 can in principle be used to construct a plug-in estimator of the DTTs. Here, we provide an informal outline of the steps. Section LABEL:sec:proofEstimation of the supplementary material considers the estimation procedure for the DTTs more formally. Introducing notations for the rescaled and recentered quantities, let ${\boldsymbol{\eta}}_{gt}^{(d)^{*}}:=\frac{{\boldsymbol{\eta}}_{gt}^{(d)}}{\left\|{\boldsymbol{\eta}}_{gt}^{(d)}\right\|}$, $\kappa_{j,gt}^{(d)^{*}}:=\frac{\kappa_{j}}{\left\|{\boldsymbol{\eta}}_{gt}^{(d)}\right\|}$, $\widetilde{\varepsilon}_{gt}^{(d)}:=\eta_{c,gt}^{(d)}+\lambda_{gt}^{(d)}\varepsilon_{gt}^{(d)}$ and $\widetilde{\varepsilon}_{gt}^{(d)^{*}}:=\frac{\widetilde{\varepsilon}_{gt}^{(d)}}{\left\|{\boldsymbol{\eta}}_{gt}^{(d)}\right\|}$ for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\}$. For simplicity, we will assume that $\eta_{c,gt}^{(d)}=\eta_{c}$ and $\lambda_{gt}^{(d)}=\lambda$ for all $(g,t)\in\{0,1\}$ and $d\in\{0,1\}$, and comment on the case where these parameters differ at the end of this remark. Given the new notation, the model in (13) can thus be reformulated as:
$Y_{gt}(d)=$ $\sum_{j=0}^{2}j\cdot\mathbb{I}\left[\kappa_{j,gt}^{(d)^{*}}<{\mathbf{X}}_{gt}^{\prime}{\boldsymbol{\eta}}^{(d)^{*}}_{gt}+\widetilde{\varepsilon}^{(d)^{*}}_{gt}\leq\kappa_{j+1,gt}^{(d)^{*}}\right].$ (22)
The rescaled slope coefficient vector ${\boldsymbol{\eta}}^{(d)^{*}}_{gt}$ in (22) can be estimated using various rank estimators available in the literature (e.g., Sherman, 1993). Subsequently, the threshold parameter $\kappa_{2,gt}^{(d)^{*}}=1/\|{\boldsymbol{\eta}}^{(d)}_{gt}\|$ can be estimated using the $\sqrt{n}_{gt}$-consistent threshold estimator for semiparametric ordered choice models proposed by Klein and Sherman (2002). This yields an estimator for the scaling factor as
$\widehat{\|\boldsymbol{\eta}^{(d)}_{gt}\|}:=1/\widehat{\kappa_{2,gt}^{(d)^{*}}}.$
Consequently, the estimator for the unscaled parameter vector becomes:
$\widehat{\boldsymbol{\eta}}^{(d)}_{gt}:=\widehat{\boldsymbol{\eta}}_{gt}^{(d)^{*}}\widehat{\|\boldsymbol{\eta}^{(d)}_{gt}\|}.$
After estimating all parameters, we recover the distributions of $\widetilde{\varepsilon}_{gt}^{(d)}$ using, e.g., a standard nonparametric kernel estimator:
$\widehat{F}_{\widetilde{\varepsilon}_{gt}^{(d)}}(v):=\widehat{\mathrm{E}}\left[\mathbb{I}\left[Y_{i}\leq 0\right]\mid\mathbf{X}_{i}^{\prime}\widehat{\boldsymbol{\eta}}_{gt}^{(d)}=-v,G_{i}=g,T_{i}=t\right].$
With estimates for all groups and time periods at hand, and under strict stationarity for the counterfactual distribution from Assumption 1, we can construct the estimated conditional DTTs for $j\in\{0,1,2\}$ and $\mathbf{x}\in\mathcal{X}_{11}$ as outlined in Proposition 6:
$\widehat{\tau}^{c}_{11}(j\mid\mathbf{x})$

### v3

@@ 4.2.1 Counterfactual Distribution
It should be emphasized that the CiC conditions solely set constraints on the change of the marginal distributions of counterfactual outcomes over time. However, these assumptions do not place any constraints on the evolution of the dependence structure between the outcome and reporting decisions. Additionally, it is noteworthy that the identification result presented in Proposition 3 requires knowledge of neither the marginal distributions nor the Copula function, thereby classifying it as semiparametric. Nevertheless, when transitioning to estimation, one has to select specific parametric forms for the Copula function $\overline{C}_{gt}(\cdot,\cdot)$ as well as the distributions $F_{-\varepsilon_{gt}}$ and $F_{-\nu_{gt}}$. This is because the identification based on the real analytic assumption is not “constructive” and does not provide a direct pathway to construct estimators.
4.2.2 Distributional Treatment Effects
@@
To identify the different DTTs, we start by setting up a model for $Y_{11}(1)$ and $R_{11}(1)$, respectively. As modeling $Y_{11}(1)$ become indispensable with misreporting, we postulate an analogous model structure for $Y_{11}(1)$ as for $Y_{11}(0)$ without misreporting. Specifically, drawing on the model for $Y_{11}^{\ast}(0)$ in (13), we define the latent consumption in the presence of treatment as:
@@
where we introduced the superscript (1) to distinguish the parameters of the potential outcome under treatment from those under control. From this point forward, we will accordingly add the superscript $(0)$ to parameters in the absence of treatment. Drawing on the model for $R_{11}^{\ast}(0)$ in (14), we define analogously:
$R_{11}^{\ast}(1)=\overline{\mathbf{Z}}_{11}^{\prime}\overline{\boldsymbol{\pi}}_{11}^{(1)}+\zeta_{11}^{(1)}\nu_{11}^{(1)},$ (19)
and adopt the following assumption:
@@
$-\varepsilon_{11}^{(1)}$ and $-\nu_{11}^{(1)}$ are strictly stationary within each group with zero median and unit variance, jointly independent of $(\mathbf{X}_{11}^{\prime},\mathbf{Z}_{11}^{\prime})^{\prime}$, and have unknown, real analytic and strictly monotonic CDFs $F_{-\varepsilon_{11}^{(1)}}$ and $F_{-\nu_{11}^{(1)}}$, respectively. Also, the copula function $\overline{C}_{11}^{(1)}(\cdot,\cdot)$ defined below is strictly monotonic and real analytic in each of its coordinates:
@@
Suppose Assumptions 6, 7, and 8 hold, and consider the model in (13)-(14) and (18)-(19). Also, suppose the CiC model from Definition 1 holds for $Y_{gt}^{*}(0)$ as well as $R_{gt}^{*}(0)$. Then, for all $j\in\{0,1,2\}$, $\mathbf{x}\in\mathcal{X}_{11}$ and $\mathbf{z}\in\mathcal{Z}_{11}$:
(i) Under Assumption 7 (i)-(a), $\tau^{c}(j)$ is identified as:
$\tau^{c}(j\mid{\mathbf{x}})=$ $\left(F_{-\varepsilon_{11}^{(1)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}-\kappa_{j+1}}{\lambda^{(1)}_{11}}\right)-F_{-\varepsilon_{11}^{(1)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}-\kappa_{j}}{\lambda^{(1)}_{11}}\right)\right)$
@@
(ii) Under Assumption 7 (i)-(b), $\tau^{r}(j)$ is identified as:
$\tau^{r}(j\mid{\mathbf{z}})=$ $\left(F_{-\nu_{11}^{(1)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(1)}_{11}-\iota_{j+1}}{\zeta^{(1)}_{11}}\right)-F_{-\nu_{11}^{(1)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(1)}_{11}-\iota_{j}}{\zeta^{(1)}_{11}}\right)\right)$
@@
Remark 1: As the identification argument presented in this subsection is not constructive, estimation may be carried out via Maximum Likelihood (ML) relying on a specific choice for the copula function. In Section S.3.2 of the supplementary material, we present corresponding ML estimators for panel as well as repeated cross-sectional data, and discuss their asymptotic properties. Specifically, as all model parameters are estimated jointly, note that consistent estimation requires both Assumption 7 (i)-(a) and (i)-(b) to hold simultaneously. For a given choice of the marginal distribution $F_{-\varepsilon}$ and $F_{-\nu}$, respectively, the ML estimators of the conditional DTTs outlined in Proposition 4 are given by:
@@
$\widehat{\tau}^{c}(j\mid\mathbf{x})$ $:=\left[F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(1)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(1)}_{11}}\right)\right]-\left[F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(0)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(0)}_{11}}\right)\right]$